OpenStudy (anonymous):

Solve each quadratic equation using factoring?? Pleasee help :) 2x^2+12=11x

OpenStudy (cgreenwade2000):

Not even a quadratic equation

OpenStudy (theviper):

is it 2x^2+12=11x

OpenStudy (anonymous):

yeah my bad

OpenStudy (anonymous):

i fixed it

OpenStudy (cgreenwade2000):

Set the equation so you get ax^2+bx+c = 0 That's step one.

OpenStudy (anonymous):

firs get everything on one side: 9x-12=0 then you can factor our a 3: so 3(3x-4)=0 divide 3 and add 4: 3x=4 x=3/4

OpenStudy (cgreenwade2000):

James, that's the wrong equation.

OpenStudy (anonymous):

@cgreenwade2000 noo its not he already simplified it

OpenStudy (theviper):

2x^2+12=11x 2x^2-11x+12=0 2x^2-8x-3x+12=0 2x(x-4)-3(x-4)=0 (2x-3)(x-4)=0

OpenStudy (anonymous):

thanks @TheViper your awesome :))) look out for my next ones

OpenStudy (cgreenwade2000):

Viper, she needs to understand how to do the problem.

OpenStudy (anonymous):

@cgreenwade2000 i do understand how to do it

OpenStudy (cgreenwade2000):

So you can factor the equation x^2+2x+1=0

OpenStudy (anonymous):

yeah i'm just too lazy to do it i'm in a rush i have other homework to do i just dont wanna sit here for hours doing a math problem

OpenStudy (cgreenwade2000):

It's good practice. And x^2+2x+1=0 is really easy to factor.